OpenStudy (anonymous):

A rectangular field measures 50 yards by 100 yards. If it takes one quart of paint to paint a 15 foot by 30 foot section, how many gallons of paint will be needed to paint the entire field

OpenStudy (anonymous):

area of the rectangular field = L * B = 50 yards * 100 yards = 5000 sq.yards one quart paints area = 15 ft x 30 ft = 450 sqft = 50sqyards paint required for 5000 sqyards = 5000/50 = 100 quarts = 25 gallons